OpenStudy (anonymous):

graph each function then identify the domain and range of the function. y=-4x^2

OpenStudy (anonymous):

hello im not sure exactly how to set this up these are quadratic functions by the way

OpenStudy (anonymous):

 The domain is (-∞,∞), while the range is (-∞,0)

OpenStudy (anonymous):

how did you get that though?

OpenStudy (anonymous):

the domain is all possible values of x. when you graph a quadratic, you will end up with a parabola

OpenStudy (anonymous):

did you get those coordinates

OpenStudy (anonymous):

did you use -2 0 and 2

OpenStudy (anonymous):

no, the graph is just a sketch of what it would look like, if it were graphed

OpenStudy (anonymous):

ok so if you were to graph it though thats how it would look like because i need to graph

OpenStudy (anonymous):

it by the coordi

OpenStudy (anonymous):

coordinates sorry my computer is going craxy now lol

OpenStudy (anonymous):

im soo sorry

OpenStudy (anonymous):

ok so how will i actually graph it?

OpenStudy (anonymous):

so, the graph of a quadratic equation will be a parabola. do you know what a parabola is?

OpenStudy (anonymous):

yes it is the u shape on the graph

OpenStudy (anonymous):

a quadratic written in standard form ax^2 + bx + c = f(x) will have vertex (-b/2a , f(-b/2a)) y = -4x^2 has: a = -4 b = 0 c = 0 so the vertex is at (0,0) Then, since a is negative, we know that the parabola will open downward.

OpenStudy (anonymous):

ok so the vertex is at 0 and the parabola is facing down when we graph

OpenStudy (anonymous):

The graph I drew was just a sketch. Here you go:

OpenStudy (anonymous):

do you have any questions?

OpenStudy (anonymous):

ok so when graphing i will take -4 and put 1in x so i get -4(1)^2 and i will solve to get -4 and i will do the same with 2 so i get -16 and i will get the graph

OpenStudy (anonymous):

example\

OpenStudy (anonymous):